# Policy Improvement using Language Feedback Models

### v3

@@ Appendix A Verbalization of visual environments
ALFWorld
SciWorld
Touchdown
Ins len $|x|$
8.8
64.7
93.4
Obs len $|v|$
23.9
239.4
284.9
Act len $|a|$
4.5
6.0
2.4
Traj len $|\tau|$
19.7
55.1
34.2
||||Act space||||
29.9
1.9k
2.1
# act |$|\{a\}|$
2.6k
2.4k
8
# ins |$|\{\tau\}|$
1.0k
1.2k
6.5k
# obs |$|\{v\}|$
18.2k
157k
34.3k
# demos
3.5k
3.6k
6.5k
@@
As shown in Figure 2, Touchdown (Chen & Mooney, 2011) is primarily a visual navigation problem. Given a set of connected Google Streetview panoramas that represent neighbourhoods in Manhattan, an agent must follow long, complex natural language instructions to navigate to the correct location. Crucial to this task of navigation are landmarks referred to by the instructions. For instance, the instruction “turn so the scaffolding is on your left and… to the next corner and turn right…” refers to the landmarks scaffolding and next corner. Prior work in verbalization use LLMs to identify landmarks (Schumann et al., 2024). In this work, we take the set of common noun-phrases in the corpus of instructions to be landmarks.
@@
First, we identify all noun-phrases using SpaCy (Honnibal & Montani, 2017). Given a visual scene, we divide the scene into 300x300 pixel non-overlapping patches. For each patch, we identify the noun-phrase with the highest cosine similarity between the noun-phrase text embedding and the image patch embedding. We use text and visual encoders from Clip (Radford et al., 2021) to extract embeddings for each modality. For patches with no aligned noun-phrase with cosine similarity greater than 0.2, we do not provide annotated a noun-phrase.
@@
It is important to note that Touchdown consists of multiple subtasks, such as finding the correct scene, stopping at the correct scene, and then finally orienting to face a hidden object in the scene. In this work, similar to Zhong et al. (2021), we only consider the first task of finding the correct scene according to the instruction. To compare our verbalization with prior work, we also evaluate a separate setting (e.g. used in Chen et al. (2019); Schumann et al. (2024)) where the agent must identify when to stop and is credited so long as it stops within one panorama of the target scene. Behaviour cloning using our verbalization technique results in 63.0% task-completion rate compared to 26.4% (Schumann et al., 2024).
@@
In Appendix A Table 6, we show statistics of verbalized environments as quantitative evidence of their challenges.
Appendix B Pseudocode for Policy Improvement using Language Feedback Models
@@
Appendix C Comparison to Dagger
@@
ALFWorld
ScienceWorld
Touchdown
Bc
62.6
45.8
57.5
ActPred
56.0
39.0
58.0
Dagger
55.2
22.5
50.2
@@
Dagger (Ross et al., 2011) is an intermediate method between Language Feedback Models and using LLMs as an expert policies for imitation learning. Specifically, in Dagger, we also use LLMs as experts to predict action. However, instead of using LLMs during each step, in Dagger we use LLMs to provide batched retroactive action prediction similar to how in Language Feedback Models we use LLMs to provide batched retroactive feedback. Here, we apply Dagger action prediction to the exact same number of examples as when we collect feedback data for LFMs. In Table 7, we compare Dagger performance to those using LLM as an expert (ActPred) and using Language Feedback Models (Lfm). We find that although Dagger is more efficient than ActPred in that it annotates synthetic examples in batch, it underperforms ActPred (and consequently Lfm) across all benchmarks.
Appendix D Quantitative and Qualitative Analyses of Learned Language Feedback
@@
GPT-4 only
Llama 2 only
both
ALFWorld
14.4%
49.3%
36.2%
ScienceWorld
10.2%
62.3%
27.5%
Touchdown
22.3%
67.3%
10.4%
@@
TP
FP
TP
FP
ALFWorld
100% (22)
0
32% (18)
68% (38)
ScienceWorld
78% (38)
22% (11)
48% (38)
52% (41)
Touchdown
81% (22)
19% (5)
39% (24)
61% (38)
@@
Step 6 Your action: go to countertop 1 Result: You arrive at loc 11. On the countertop 1, you see a candle 2, a candle 1, a soapbar 2…
Step 11 Your action: open cabinet 1 Result: You open the cabinet 1. The cabinet 1 is open. In it, you see a candle 3, and a toiletpaper 1.
Step 17 Your action: open cabinet 1 Result: You open the cabinet 1. The cabinet 1 is open. In it, you see a candle 3, and a toiletpaper 1.
@@
Step 11 Your action: open drawer 3 Result: You open the drawer 3. The drawer 3 is open. In it, you see nothing.
Step 14 Your action: open drawer 4 Result: You open the drawer 4. The drawer 4 is open. In it, you see nothing.
Step 15 Your action: close drawer 4 Result: You close the drawer 4.
Step 19 Your action: go to drawer 6 Result: You arrive at loc 27. The drawer 6 is closed.
Step 20 Your action: open drawer 6 Result: You open the drawer 6. The drawer 6 is open. In it, you see nothing.
@@
Step 38 Your action: move aluminum foil to yellow box Result: You move the aluminum foil to the yellow box…
@@
Step 13 Your action: pick up unknown substance Result: You move the unknown substance N to the inventory…
Step 15 Your action: move unknown substance to blue box Result: You move the unknown substance N to the blue box…
@@
Step 1 Your action: slightly to your left Result: behind you, you see: parked, white box trucks , two wider sidewalks , this narrow two lane road , the right sidewalk buildings . behind you, sightly to your left, you see: , three air-conditioners , three awning , a smaller yellow taxi . to your left, you see: , the theater awning , a yellow cab car , the second purple awning . slightly to your left, you see: , a white-capped hydrant , ornate gray balconies , the purple wayfair truck . straight ahead, you see: , a median strip , some tall, brick buildings , parked, white box trucks , a blue bus lane sign , the right sidewalk buildings . slightly to your right, you see: , this nearest intersection , a light colored stone building , a blue and red bus stop sign to your right, you see: , separate building entrances , the half-moon curb , the company’s name . behind you, slightly to your right, you see: , , then a storefront , ornate gray balconies , the first little weird intersection.
@@
Step 8 Your action: straight ahead Result: behind you, you see: , a brown storefront , surface streets , the right sidewalk buildings , a parked black box truck , some unremarkable brick buildings . behind you, sightly to your left, you see: , then a storefront , a white/grey van , a large, blocky, gray building . to your left, you see: , a large white store sign , a construction vehicle , a long gray and white 5 story building slightly to your left, you see: , some tall, brick buildings , fedex van , a white/grey van . straight ahead, you see: , a large red brick apartment building , an orange and white traffic object , 3rd and 4th intersections , a small blue car , the right sidewalk buildings , the parked yellow suv taxi . slightly to your right, you see: , construction awning , a large white fedex truck , a white/grey van . to your right, you see: , this van stop , the building edge , a large white fedex truck . behind you, slightly to your right, you see: , a green-topped scaffolding , a large white fedex truck , a parked black box truck.
Comparison of GPT-4 to Llama 2 70B
Appendix E GPU Usage

### v6

@@ Appendix D Verbalization of visual environments
ALFWorld SciWorld Touchdown
Ins len $|x|$ 8.8 64.7 93.4
Obs len $|v|$ 23.9 239.4 284.9
Act len $|a|$ 4.5 6.0 2.4
Traj len $|\tau|$ 19.7 55.1 34.2
||||Act space|||| 29.9 1.9k 2.1
# act |$|\{a\}|$ 2.6k 2.4k 8
# ins |$|\{\tau\}|$ 1.0k 1.2k 6.5k
# obs |$|\{v\}|$ 18.2k 157k 34.3k
# demos 3.5k 3.6k 6.5k
@@
As shown in Figure 2, Touchdown [11] is primarily a visual navigation problem. Given a set of connected Google Streetview panoramas that represent neighbourhoods in Manhattan, an agent must follow long, complex natural language instructions to navigate to the correct location. Crucial to this task of navigation are landmarks referred to by the instructions. For instance, the instruction “turn so the scaffolding is on your left and… to the next corner and turn right…” refers to the landmarks scaffolding and next corner. Prior work in verbalization use LLMs to identify landmarks [34]. In this work, we take the set of common noun-phrases in the corpus of instructions to be landmarks.
@@
First, we identify all noun-phrases using SpaCy [20]. Given a visual scene, we divide the scene into 300x300 pixel non-overlapping patches. For each patch, we identify the noun-phrase with the highest cosine similarity between the noun-phrase text embedding and the image patch embedding. We use text and visual encoders from Clip [30] to extract embeddings for each modality. For patches with no aligned noun-phrase with cosine similarity greater than 0.2, we do not provide annotated a noun-phrase.
@@
It is important to note that Touchdown consists of multiple subtasks, such as finding the correct scene, stopping at the correct scene, and then finally orienting to face a hidden object in the scene. In this work, similar to Zhong et al. [49], we only consider the first task of finding the correct scene according to the instruction. To compare our verbalization with prior work, we also evaluate a separate setting (e.g. used in Chen et al. [12], Schumann et al. [34]) where the agent must identify when to stop and is credited so long as it stops within one panorama of the target scene. Behaviour cloning using our verbalization technique results in 63.0% task-completion rate.
@@
In Appendix D Table 6, we show statistics of verbalized environments as quantitative evidence of their challenges.
Appendix E Pseudocode for Policy Improvement using Language Feedback Models
@@
Appendix F Comparison to Dagger
@@
ALFWorld ScienceWorld Touchdown
Bc 62.6 45.8 57.5
ActPred 56.0 39.0 58.0
Dagger 55.2 22.5 50.2
@@
Dagger [31] is an intermediate method between Language Feedback Models and using LLMs as an expert policies for imitation learning. Specifically, in Dagger, we also use LLMs as experts to predict action. However, instead of using LLMs during each step, in Dagger we use LLMs to provide batched retroactive action prediction similar to how in Language Feedback Models we use LLMs to provide batched retroactive feedback. Here, we apply Dagger action prediction to the exact same number of examples as when we collect feedback data for LFMs. In Table 7, we compare Dagger performance to those using LLM as an expert (ActPred) and using Language Feedback Models (Lfm). We find that although Dagger is more efficient than ActPred in that it annotates synthetic examples in batch, it underperforms ActPred (and consequently Lfm) across all benchmarks.
Appendix G Quantitative and Qualitative Analyses of Learned Language Feedback
@@
GPT-4 only Llama 2 only both
ALFWorld 14.4% 49.3% 36.2%
ScienceWorld 10.2% 62.3% 27.5%
Touchdown 22.3% 67.3% 10.4%
@@
TP FP TP FP
ALFWorld 100% (22) 0 32% (18) 68% (38)
ScienceWorld 78% (38) 22% (11) 48% (38) 52% (41)
Touchdown 81% (22) 19% (5) 39% (24) 61% (38)
Comparison of GPT-4 to Llama 2 70B
Appendix H GPU Usage
Appendix I Correlation between LLM/LFM feedback and true state values
To investigate whether LLM feedback is correlated with true state values, we obtained partial rollouts for environments in ALFWorld and asked GPT4 to score from 1-5 whether the partial rollouts are on the right track to solving the task. We then ran a planner (with full observability) to complete these partial rollouts in order to obtain ground truth optimal values. With no training, the LLM’s predicted score has strong correlation (0.61 Pearson) with the optimal values. This means that GPT as a feedback model has strong zero-shot generalization on environments it was not trained on.
In addition, we regressed a FLAN-T5 model to estimate state values from language feedback using states from a random policy, then evaluated its predictions against true state values on mixed policies where an expert is select $p$ fraction of the time (and a random policy is used other times). When we uniformly sample states from the random policy, for $p=0,0.2,0.4,0.6,0.8$, the regression model achieves Pearson correlations of $0.3-0.4$. When we subsample training states evenly across values, we achieve Pearson $0.65-0.75$. This shows that state coverage, for which state value coverage is a proxy, is indeed important. Conversely, when we use mixed policies for training as well, we obtain $0.78-0.85$ Pearson. This shows that optimality is indeed important.
@@ Appendix I Correlation between LLM/LFM feedback and true state values
Step 6 Your action: go to countertop 1 Result: You arrive at loc 11. On the countertop 1, you see a candle 2…
Step 11 Your action: open cabinet 1 Result: You open the cabinet 1. The cabinet 1 is open. In it, you…
Step 17 Your action: open cabinet 1 Result: You open the cabinet 1. The cabinet 1 is open. In it, you see…
@@
Step 11 Your action: open drawer 3 Result: You open the drawer 3. The drawer 3 is open. In it, you…
Step 14 Your action: open drawer 4 Result: You open the drawer 4. The drawer 4 is open. In it, you…
@@
Step 38 Your action: move aluminum foil to yellow box Result: You move the aluminum foil to the yellow…
@@
Step 13 Your action: pick up unknown substance Result: You move the unknown substance N to the…
Step 15 Your action: move unknown substance to blue box Result: You move the unknown substance N to…
@@
Step 1 Your action: slightly to your left Result: behind you, you see: parked, white box trucks , two wider sidewalks , this narrow two lane road , the right sidewalk buildings . behind you, sightly to your left, you see: , three air-conditioners , three awning , a smaller yellow taxi . to your left, you see: , the theater awning , a yellow cab car , the second purple awning . slightly to your left, you see: , a white-capped hydrant , ornate gray balconies , the purple wayfair truck . straight ahead, you see: , a median strip , some tall, brick buildings , parked, white box trucks , a blue bus lane sign , the right sidewalk buildings . …
@@
Step 8 Your action: straight ahead Result: behind you, you see: , a brown storefront , surface streets , the right sidewalk buildings , a parked black box truck , some unremarkable brick buildings . behind you, sightly to your left, you see: , then a storefront , a white/grey van , a large, blocky, gray building . to your left, you see: , a large white store sign , a construction vehicle , a long gray and white 5 story building slightly to your left, you see: , some tall, brick buildings , fedex van , a white/grey van . straight ahead, you see: , a large red brick apartment building , an orange and white traffic object , 3rd and 4th intersections , a small blue car , the right sidewalk buildings , the parked yellow suv taxi . slightly to your right, you see: …
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: we outline a new method for policy improvement via Language Feedback Model. In the abstract and in the introduction, we claim that this method improves policy performance. Our experiments show that this method improves base policy performance on 3 distinct benchmarks. Furthermore, our experiments show that this technique outperforms alternative methods such as using LLMs as zeroshot policy, using LLMs to provide action labels, and using LLMs to provide Dagger-style annotations.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We discuss the limitations of this work in Appendix A.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: This work does not include theoretical results.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We outline experimental details in this paper, including steps and prompts required to produce LLM feedback to train LFM. Furthermore, we outline hyperparameters and compute resources used to train feedback models and policies. Finally, we release our code and checkpoints for reproducing experimental results.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We open source data and code required to reproduce our experiments. This release also contains documentation on how to run the code to reproduce data and re-train models described in this work. At this time, the release is anonymized at the link described in the introduction.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: Our experiment section describe hyperparameters used to train models and collection feedback data. They are also contained in the source code release.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: We provide standard deviation of task completion rates in our main results table in Table 3 across 3 random seeds.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We describe compute usage in Appendix H.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: we have reviewed the NeurIPS Code of Ethics and made sure that this work does not deviate from the Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: We discuss the broader impact of this work in Appendix B.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: This work does not release data or models that have a high risk of misuse. We discuss limitations and impact of released models in Appendix A and Appendix B.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We detail license and terms of usage of models released for this work as well as assets used to develop this work in Appendix C.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: This work does not release new assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: This work does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: This work does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.